# The MUSE eXtremely Deep Field: Detections of circumgalactic Siāii*Ā emission at ā ā thanks: Based on observations made with ESO telescope at the La Silla Paranal Observatory under the large program 1101.A-0127

### v1

@@ 4.1 Testing photon conservation
Shapley etĀ al. (2003) suggest that the origin of the Siāii*Ā is continuum pumping (see Section 1 for other origins). For the Siāii* line, Si+ in neutral clouds makes a Siāii resonant absorption line, and then a fraction of Si+ deexcites to the fine-structure level, emitting SiāiiĀ fluorescence photons at 1533Ā Ć Ā (see Figure 1). As the SiāiiĀ is a resonant line, the Siāii*Ā line can be regarded as an escape channel of photons from resonant scattering. If dust extinction is negligible and the escape is isotropic, then the equivalent width (EW) of Siāii*Ā emission and that of the corresponding resonant absorption should be comparable because of photon conservation. However, it is possible for the absorption and emission to have different EWs for one given direction of observations even if the origin is continuum pumping without dust, because of the complex gas distribution in front of stars in galaxies (e.g., Prochaska etĀ al. 2011b; Carr etĀ al. 2018). Simulations predict that absorption lines depend sensitively on the direction of observations (Mauerhofer etĀ al. 2021; Gazagnes etĀ al. 2023). Below, we assume an isotropic case.
@@
To investigate the scenario of continuum pumping, we use the isolated, high-SN object MID (see Section 3.2) and test photon conservation between SiāiiĀ absorption and associated Siāii*Ā emission, which is commonly assessed with the EWs of these lines (e.g., Prochaska etĀ al. 2011a; Wang etĀ al. 2020). First, we check the flux curve of growth (CoG) for the continuum and the emission line to determine the spatial extent. Second, we measure the EW curve of growth for Siāii*Ā emission and compare it with the absorption EW. We measure EWs from the MUSE data (without HST), which makes the EW comparison fair. Ideally, we should measure the curve of growth for the absorption. However, due to the limited depth and spatial resolution of our data, we have to assume that the absorption follows the same spatial profile as that of the continuum. The left panel of Figure 5 shows normalized fluxes within growing apertures for continuum (black) and Siāii* (blue). Fluxes are measured from the Siāii* NB image and the corresponding UV continuum BB image (Section 3.1) with PHOTUTILS and the neighboring object mask with corrections of masked areas. On a galaxy-scale of R, the aperture includes about 80% of the continuum but only 50% of the Siāii*. An R aperture captures most of the Siāii*Ā flux.
@@
The right panel shows rest-frame EWs for Siāii* emission within growing apertures (blue), compared with the EW for Siāii absorption measured at R, EWabs(Siāii)Ā Ć Ā (red). Rest-frame EWs are measured in spectra extracted from the original minicube (to which the neighboring object mask is applied) with growing apertures. The continuum flux density is measured by the median filtering each spectrum from 50-slice shorter wavelength (62.5Ć ) than the absorption to 50-slice longer wavelength than the emission. Absorption and emission fluxes are measured as differences from the continuum. The EW for Siāii* is EWem(Siāii*)Ā Ć Ā at R, which is smaller than EWabs(Siāii)Ā Ć . EWem(Siāii*) increases with R and reaches EWem(Siāii*)Ā Ć Ā at R, which is consistent with EWabs(Siāii) at R within the 1 uncertainties (see Figure 6 for the spectra), although the EWs have large measurement errors. Therefore, the absorption EW on the galaxy scale and the fluorescent emission EW on the CGM scale are consistent with each other. This is compatible with photon conservation on the CGM scale, albeit with large uncertainties. It implies that the origin of extended Siāii*Ā emission can be the continuum pumping as predicted under the assumption of isotropic escape and no dust. In that case, either a small amount of dust, a low optical depth, or a combination of both are implied. Unfortunately, data for the other four sources with individual Siāii*Ā halos are either noisy or have accompanying galaxies inside their 5ā5ā minicubes, which prevents us from obtaining a firm conclusion about the mechanisms for the four sources. The trend of increasing EW of Siāii*Ā emission with radius beyond the galaxy scale is predicted by simulations (Mauerhofer etĀ al. 2021), which are compared with slit spectroscopy of low-redshift galaxies in Gazagnes etĀ al. (2023, see also ).
@@ 4.2 Comparisons with zoom simulations
We compare our stacked Siāii* profile with those from the cosmological zoom-in simulations of Mauerhofer etĀ al. (2021), which are used in Gazagnes etĀ al. (2023) and Blaizot etĀ al. (2023). Mauerhofer etĀ al. (2021) present three snapshots containing a relatively small galaxy (such as MUSE LAEs) at , , and . The stellar mass for the galaxy is Ā Mā, and the star formation rate ranges from to 5Ā MāĀ yr-1 for the three snapshots. Siāii*photons in this simulations originate from continuum pumping. We provide mock MUSE cubes for observations from 12 directions for each snapshot in the rest frame. This original mock cube has a size of pixels corresponding to Ā Ć Ā (from 1521.83Ā Ć Ā to 1538.42Ā Ć ) with a spaxel scale of /pix. Then, to match the size of observed and simulated galaxies, we measure the median size of the simulated galaxy among the 36 directions. A UV-continuum BB image is created from each median-filtered original cube with a window from 1528.5Ā Ć Ā to 1531.5Ā Ć Ā (the continuum between SiāiiĀ absorption and Siāii*Ā emission). We define the galaxy center by identifying the brightest pixel, as in Gazagnes etĀ al. (2023), for each UV continuum BB image after HST PSF convolution (F606W in Rafelski etĀ al. 2015). The half-light radius (R50) is measured around the galaxy center with the original mock cube (without HST PSF convolution) using PHOTUTILS. The median R50 for the simulated galaxy is (3.5 pix). This is 1.7 times smaller than the median R50 for the 14 stacked galaxies (R50= with F606W after PSF correction). Since we do not have a simulated galaxy that is large enough in Mauerhofer etĀ al. (2021), we manually rescale the original mock cube by a factor of 1.7 by interpreting the spaxel scale as /pix, which is referred to as an intrinsic mock cube below. The intrinsic mock cubes are cut out around the galaxy center defined above and are 5 arcsec each. The 5ā cubes are convolved with the MUSE PSF and line spread function (LSF) described in Bacon etĀ al. (2023) at the median redshift of the stacked sample (). Then the convolved mock cubes are rebinned to match the MUSE cubes in the rest frame. We apply similar analyses to the 36 mock MUSE cubes as used for the observational data (see Section 3). A slight difference from the method for the MUSE data is continuum subtraction. Continuum cubes used for continuum subtraction are created using median filtering from 1528.5Ā Ć Ā to 1531.5Ā Ć ā which are also used for the UV-continuum BB images. The 1 uncertainties for stacked mock data are derived from 15.87 to 84.12 percentiles of SB profiles among individual mocks. We would like to note that the simulations assume cosmological parameters of km s-1 Mpc-1 and , which should not have a significant impact on the results below. We convert distances in the unit of arcsec for the simulated data to those in the unit of kpc at with the Planck 2018 cosmological model (Planck Collaboration etĀ al. 2020) as the galaxy sizes are matched and MUSE PSF are convolved in the unit of arcsec.
@@
The left panel of Figure 7 shows a comparison of the observed and simulated mean-stacked SB profiles. The simulated UV continuum (gray dotted line) is well matched with the observations (black solid line), thanks to our manual 1.7 times expansion. The simulated Siāii* profile (green dotted line) is more extended than the simulated UV continuum (gray dotted line) and consistent with the observed Siāii* (blue solid line), though simulated Siāii*Ā halo is slightly less extended than observed within uncertainties.
@@
The mean profiles can be biased toward bright halo sources, so we check whether the mean-stacked profiles are consistent with the median-stacked profiles for the Siāii*Ā and the continuum (green and gray crosses, respectively) in the right panel of Figure 7. We confirm that they are consistent with each other, as for the MUSE observations of Siāii*. The right panel also compares simulated mean-stacked SB profiles with PSF and LSF convolution versus without convolution, indicated by dotted and solid lines, respectively. The simulated galaxy indeed has significantly more extended Siāii* than the UV continuum in the intrinsic mock cubes, but the difference is mostly hidden by the MUSE PSF convolution555We check individual mocks and find that the shapes of the individual SB profiles (green dotted line) are fully dominated by the PSF..
@@
We also check EW CoG of Siāii* emission and Siāii absorption lines for the simulations as shown in the left panel of Figure 8. The EWs of Siāii* emission and Siāii absorption are consistent at a CGM scale for both intrinsic cubes and PSF-convolved cubes, which are also seen on the mock spectra (see the middle and right panels of Figure 8). It means that the photon conservation works for the simulations with continuum pumping in the sense that the scattered photons are not more attenuated by dust than that stellar continuum. This is consistent with our observations for MID=1141.
@@
From these comparisons, we conclude that our simulations with continuum pumping scenario can reproduce the observations. This lends extra support for the interpretation of Siāii*Ā emission as a signature of continuum pumping. We note that the selection bias with the UV magnitude cut may have an impact on the observed result and this comparison. The mock observations are re-scaled by a factor of to correct for the fact that the simulated galaxy is smaller than the observed ones (see above). The simulated galaxy is fainter than the observed one, though the range of absolute magnitudes overlaps ( from -18 to -19). It would be interesting to test the selection bias with a larger and brighter simulated galaxy, or several, maybe at a slightly lower redshift.
@@ 4.3 Possible extension of the sample
We discuss possible strategies to extend the sample with existing data in this section. There are a few ways to increase the sample size: (1) sources at , which have other nebular line detections such as the Oāiii]1661, 1666 doublet and Heāii in the same catalog and the same field (with the same integration time), (2) sources located in shallower fields such as mosaicĀ field (99-pointing with 10-hour integration, in the same catalog of Bacon etĀ al. 2023) and MUSE-WIDE field (1-hour integration Herenz etĀ al. 2017; Urrutia etĀ al. 2019; Kerutt etĀ al. 2022), (3) UV-bright sources at the same redshifts in other fields with similarly long integration times ( hours) in MUSE archival data (e.g., Fossati etĀ al. 2019; Lofthouse etĀ al. 2020). The samples (1) and (2) are tested with the MUSE DR2 catalog (not including MUSE WIDE) in the same manner. We find that these less restrictive selection criteria do not help to increase the number of detections of Siāii*Ā halos. Therefore, (3) using deep MUSE archival data for UV-bright galaxies with the same selection criteria would be the best way to extend the sample in the future.
In addition, (4) gravitationally lensed sources observed with MUSE (e.g., Richard etĀ al. 2021; Claeyssens etĀ al. 2022, Claeyssens et al. 2024 in prep.), would be useful to investigate more compact Siāii*Ā halos in the future (see the right panel of Figure 7), which would be hidden by the PSF without the power of the magnification. We would like to note that they are not included in this pilot study as magnification can add complications and uncertainties. To investigate Siāii*Ā halos at higher redshifts (ā), whose Siāii*Ā lines are covered with MUSE, we could use (5) sources with systemic redshifts measured by nebular lines with James Webb Space Telescope (JWST), and (6) MUSE LAEs, whose systemic redshifts can be estimated by empirical relationships based on the Ly line (Verhamme etĀ al. 2018). We do not include these sources in this pilot study, considering uncertainties in wavelength calibration (e.g., Maseda etĀ al. 2023; DāEugenio etĀ al. 2024; Meyer etĀ al. 2024) and more severe cosmic dimming effects. It is future work for us to investigate higher- sources.
@@ 5 Conclusions
To study the spatial distribution of the metal-enriched cool CGM at cosmic noon in emission, we search for extended Siāii*Ā emission (fluorescent lines) using the MUSE HUDF data and catalog with 30-140 hour integration times. We construct a sample of 39 galaxies with systemic redshifts at ā, at which the [Cāiii]1907, Cāiii]1909 doublet nebular lines and at least one of the Siāii*1265, 1309, 1533 lines are redshifted into the MUSE wavelength range. Our major results are summarized as follows.
@@
Five individual Siāii* halos are confirmed to be present from statistical tests with surface brightness profiles. These are the first detections of extended Siāii*Ā emission.
@@
We stack images of Siāii*1533 line for a subsample of 13 UV-bright galaxies. We confirm the presence of stacked Siāii* halos, which may imply that metal-enriched CGM is common for UV-bright galaxies. We also stack images of Siāii*1265, 1309, 1533 for 14 UV-bright galaxies and detect a Siāii*Ā halo. If we stack fainter galaxies or remove the five individually-detected halos from stacking, we get nondetections of Siāii*Ā halos. We need a larger sample to draw a firm conclusion.
@@
We test the presence of Siāii*Ā halos in zoom-in simulations, which account only for continuum pumping. After re-scaling the mock observations to correct for the fact that the simulated galaxy is smaller than the observed one, we find that the simulated halos are consistent with the observed stacked halo within uncertainties. This lends extra support for the interpretation of Siāii*Ā emission as a signature of continuum pumping.
@@
KCWI or BlueMUSE (Richard etĀ al. 2019) will allow us to investigate HI Ly halos for Siāii*Ā halos detected with MUSE. BlueMUSE will also make it possible to individually detect Siāii*Ā halos of more diverse sources and to study their statistical properties thanks to a more moderate cosmic dimming effect at (by a factor of three, compared to ). Such larger samples would be useful to improve the stacking experiments (see also Section 4.3) and to study spectral variations at different radii from the centers of galaxies as done for Ly (Wisotzki etĀ al. 2018; Guo etĀ al. 2023c, b). Such studies will be advanced more with better sensitivities and higher spatial resolutios achieved by extremely large telescopes such as Thirty-meter telescope (TMT), one of whose key science cases is CGM mapping with WFOS (Skidmore etĀ al. 2015), as well as European-Extremely large telescope (E-ELT).

### v2

@@ 4.1 Testing photon conservation
Shapley et al. (2003) suggest that the origin of the Si ii* is continuum pumping (see Section 1 for other origins). For the Si ii* line, Si+ in neutral clouds makes a Si ii resonant absorption line, and then a fraction of Si+ deexcites to the fine-structure level, emitting Si ii fluorescence photons at 1533 Å (see Fig. 1). As the Si ii is a resonant line, the Si ii* line can be regarded as an escape channel of photons from resonant scattering. When dust opacity is high in the ISM and the CGM, dust attenuation for Si ii* can be greater than that for the UV continuum, depending on the number of scatterings. if additional dust attenuation of Si ii* is negligible and the escape is isotropic, then the equivalent width (EW) of Si ii* emission and that of the corresponding resonant absorption should be comparable because of photon conservation.
However, it is possible for the absorption and emission to have different EWs for one given direction of observations even if the origin is continuum pumping without dust, because of the complex gas distribution in front of stars in galaxies (e.g., Prochaska et al. 2011b; Carr et al. 2018). Simulations predict that absorption lines depend sensitively on the direction of observations (Mauerhofer et al. 2021; Gazagnes et al. 2023). Below, we assume an isotropic case.
@@
To investigate the scenario of continuum pumping, we use the isolated, high-SN object MID (see Section 3.2) and test photon conservation between Si ii absorption and associated Si ii* emission, which is commonly assessed with the EWs of these lines (e.g., Prochaska et al. 2011a; Wang et al. 2020). First, we check the flux curve of growth (CoG) for the continuum and the emission line to determine the spatial extent. Second, we measure the EW curve of growth for Si ii* emission and compare it with the absorption EW. We measure EWs from the MUSE data (without HST), which makes the EW comparison fair. Ideally, we should measure the curve of growth for the absorption. However, due to the limited depth and spatial resolution of our data, we have to assume that the absorption follows the same spatial profile as that of the continuum. The left panel of Fig. 5 shows normalized fluxes within growing apertures for continuum (black) and Si ii* (blue). Fluxes are measured from the Si ii* NB image and the corresponding UV continuum BB image (Section 3.1) with PHOTUTILS and the neighboring object mask with corrections of masked areas. On a galaxy-scale of R, the aperture includes about 80% of the continuum but only 50% of the Si ii*. An R aperture captures most of the Si ii* flux.
@@
The right panel shows rest-frame EWs for Si ii* emission within growing apertures (blue), compared with the EW for Si ii absorption measured at R, EWabs(Si ii ) Å (red). Rest-frame EWs are measured in spectra extracted from the original minicube (to which the neighboring object mask is applied) with growing apertures. The continuum flux density is measured by the median filtering each spectrum from 50-slice shorter wavelength (62.5Å) than the absorption to 50-slice longer wavelength than the emission. Absorption and emission fluxes are measured as differences from the continuum. The EW for Si ii* is EWem(Si ii* ) Å at R, which is smaller than EWabs(Si ii ) Å. EWem(Si ii* ) increases with R and reaches EWem(Si ii* ) Å at R, which is consistent with EWabs(Si ii ) at R within the 1 uncertainties (see Fig. 6 for the spectra), although the EWs have large measurement errors. Therefore, the absorption EW on the galaxy scale and the fluorescent emission EW on the CGM scale are consistent with each other. This is compatible with photon conservation on the CGM scale, albeit with large uncertainties. It implies that the origin of extended Si ii* emission can be the continuum pumping as predicted under the assumption of isotropic escape and no dust. In that case, either a small amount of dust, a low optical depth, or a combination of both are implied. Unfortunately, data for the other four sources with individual Si ii* halos are either noisy or have accompanying galaxies inside their 5”5” minicubes, which prevents us from obtaining a firm conclusion about the mechanisms for the four sources. The trend of increasing EW of Si ii* emission with radius beyond the galaxy scale is predicted by simulations (Mauerhofer et al. 2021), which are compared with slit spectroscopy of low-redshift galaxies in Gazagnes et al. (2023, see also ).
@@ 4.2 Comparisons with zoom simulations
We compare our stacked Si ii* profile with those from the cosmological zoom-in simulations of Mauerhofer et al. (2021), which are used in Gazagnes et al. (2023) and Blaizot et al. (2023). Mauerhofer et al. (2021) present three snapshots containing a relatively small galaxy (such as MUSE LAEs) at , , and . The stellar mass for the simulated galaxy is M⊙, which is similar to the median value of those for the 14 galaxies used in the stacking analysis, M⊙ (ranging – M⊙, Bacon et al. 2023, see Fig. 9). The SFRs of the simulated galaxy for the three snapshots are M⊙ yr-1, M⊙ yr-1, and M⊙ yr-1, which are similar to the median value of the observed galaxies, 3.4 M⊙ yr-1 (ranging – M⊙ yr-1, Bacon et al. 2023, see Fig. 9). Si ii*photons in this simulations originate from continuum pumping. We provide mock MUSE cubes for observations from 12 directions for each snapshot in the rest frame. This original mock cube has a size of pixels corresponding to Å (from 1521.83 Å to 1538.42 Å) with a spaxel scale of /pix. Then, to match the size of observed and simulated galaxies, we measure the median size of the simulated galaxy among the 36 directions. A UV-continuum BB image is created from each median-filtered original cube with a window from 1528.5 Å to 1531.5 Å (the continuum between Si ii absorption and Si ii* emission). We define the galaxy center by identifying the brightest pixel, as in Gazagnes et al. (2023), for each UV continuum BB image after HST PSF convolution (F606W in Rafelski et al. 2015). The half-light radius (R50) is measured around the galaxy center with the original mock cube (without HST PSF convolution) using PHOTUTILS. The median R50 for the simulated galaxy is (3.5 pix). This is 1.7 times smaller than the median R50 for the 14 stacked galaxies (R50= with F606W after PSF correction). Since we do not have a simulated galaxy that is large enough in Mauerhofer et al. (2021), we manually rescale the original mock cube by a factor of 1.7 by interpreting the spaxel scale as /pix, which is referred to as an intrinsic mock cube below. The intrinsic mock cubes are cut out around the galaxy center defined above and are 5 arcsec each. The 5” cubes are convolved with the MUSE PSF and line spread function (LSF) described in Bacon et al. (2023) at the median redshift of the stacked sample (). Then the convolved mock cubes are rebinned to match the MUSE cubes in the rest frame. We apply similar analyses to the 36 mock MUSE cubes as used for the observational data (see Section 3). A slight difference from the method for the MUSE data is continuum subtraction. Continuum cubes used for continuum subtraction are created using median filtering from 1528.5 Å to 1531.5 Å which are also used for the UV-continuum BB images. The 1 uncertainties for stacked mock data are derived from 15.87 to 84.12 percentiles of SB profiles among individual mocks. We would like to note that the simulations assume cosmological parameters of km s-1 Mpc-1 and , which should not have a significant impact on the results below. We convert distances in the unit of arcsec for the simulated data to those in the unit of kpc at with the Planck 2018 cosmological model (Planck Collaboration et al. 2020) as the galaxy sizes are matched and MUSE PSF are convolved in the unit of arcsec.
@@
The left panel of Fig. 7 shows a comparison of the observed and simulated mean-stacked SB profiles. The simulated UV continuum (gray dotted line) is well matched with the observations (black solid line), thanks to our manual 1.7 times expansion. The simulated Si ii* profile (green dotted line) is more extended than the simulated UV continuum (gray dotted line) and consistent with the observed Si ii* (blue solid line), though simulated Si ii* halo is slightly less extended than observed within uncertainties.
@@
The mean profiles can be biased toward bright halo sources, so we check whether the mean-stacked profiles are consistent with the median-stacked profiles for the Si ii* and the continuum (green and gray crosses, respectively) in the right panel of Fig. 7. We confirm that they are consistent with each other, as for the MUSE observations of Si ii* . The right panel also compares simulated mean-stacked SB profiles with PSF and LSF convolution versus without convolution, indicated by dotted and solid lines, respectively. The simulated galaxy indeed has significantly more extended Si ii* than the UV continuum in the intrinsic mock cubes, but the difference is mostly hidden by the MUSE PSF convolution777We check individual mocks and find that the shapes of the individual SB profiles (green dotted line) are fully dominated by the PSF..
@@
We also check EW CoG of Si ii* emission and Si ii absorption lines for the simulations as shown in the left panel of Fig. 8. The EWs of Si ii* emission and Si ii absorption are consistent at a CGM scale for both intrinsic cubes and PSF-convolved cubes, which are also seen on the mock spectra (see the middle and right panels of Fig. 8). It means that the photon conservation works for the simulations with continuum pumping in the sense that the scattered photons are not more attenuated by dust than that stellar continuum. This is consistent with our observations for MID=1141.
@@
From these comparisons, we conclude that our simulations with continuum pumping scenario can reproduce the observations. This lends extra support for the interpretation of Si ii* emission as a signature of continuum pumping. We note that the selection bias with the UV magnitude cut may have an impact on the observed result and this comparison. The mock observations are re-scaled by a factor of to correct for the fact that the simulated galaxy is smaller than the observed ones (see above). The simulated galaxy is fainter than the observed one, though the range of absolute magnitudes overlaps ( from -18 to -19). It would be interesting to test the selection bias with a larger and brighter simulated galaxy, or several, maybe at a slightly lower redshift.
@@ 4.3 Possible extension of the sample
We discuss possible strategies to extend the sample with existing data in this section. There are a few ways to increase the sample size: (1) sources at , which have other nebular line detections such as the O iii] 1661, 1666 doublet and He ii in the same catalog and the same field (with the same integration time), (2) sources located in shallower fields such as mosaic field (99-pointing with 10-hour integration, in the same catalog of Bacon et al. 2023) and MUSE-WIDE field (1-hour integration Herenz et al. 2017; Urrutia et al. 2019; Kerutt et al. 2022), (3) UV-bright sources at the same redshifts in other fields with similarly long integration times ( hours) in MUSE archival data (e.g., Fossati et al. 2019; Lofthouse et al. 2020). The samples (1) and (2) are tested with the MUSE DR2 catalog (not including MUSE WIDE) in the same manner. We find that these less restrictive selection criteria do not help to increase the number of detections of Si ii* halos. Therefore, (3) using deep MUSE archival data for UV-bright galaxies with the same selection criteria would be the best way to extend the sample in the future.
In addition, (4) gravitationally lensed sources observed with MUSE (e.g., Richard et al. 2021; Claeyssens et al. 2022, Claeyssens et al. in prep.), would be useful to investigate more compact Si ii* halos in the future (see the right panel of Fig. 7), which would be hidden by the PSF without the power of the magnification. We would like to note that they are not included in this pilot study as magnification can add complications and uncertainties. To investigate Si ii* halos at higher redshifts (–), whose Si ii* lines are covered with MUSE, we could use (5) sources with systemic redshifts measured by nebular lines with JWST, and (6) MUSE LAEs, whose systemic redshifts can be estimated by empirical relationships based on the Ly line (Verhamme et al. 2018). We do not include these sources in this pilot study, considering uncertainties in wavelength calibration (e.g., Maseda et al. 2023; D’Eugenio et al. 2024; Meyer et al. 2024) and more severe cosmic dimming effects. It is future work for us to investigate higher- sources.
@@ 5 Conclusions
To study the spatial distribution of the metal-enriched cool CGM at cosmic noon in emission, we search for extended Si ii* emission (fluorescent lines) using the MUSE HUDF data and catalog with 30-140 hour integration times. We construct a sample of 39 galaxies with systemic redshifts at –, at which the [C iii]1907, C iii]1909 doublet nebular lines and at least one of the Si ii* 1265, 1309, 1533 lines are redshifted into the MUSE wavelength range. Our major results are summarized as follows.
@@
Five individual Si ii* halos are confirmed to be present from statistical tests with surface brightness profiles. These are the first detections of extended Si ii* emission.
@@
We stack images of Si ii* 1533 line for a subsample of 13 UV-bright galaxies. We confirm the presence of stacked Si ii* halos, which may imply that metal-enriched CGM is common for UV-bright galaxies. We also stack images of Si ii* 1265, 1309, 1533 for 14 UV-bright galaxies and detect a Si ii* halo. If we stack fainter galaxies or remove the five individually-detected halos from stacking, we get nondetections of Si ii* halos. We need a larger sample to draw a firm conclusion.
@@
We test the presence of Si ii* halos in zoom-in simulations, which account only for continuum pumping. After re-scaling the mock observations to correct for the fact that the simulated galaxy is smaller than the observed one, we find that the simulated halos are consistent with the observed stacked halo within uncertainties. This trends extra support for the interpretation of Si ii* emission as a signature of continuum pumping.
@@
KCWI or BlueMUSE (Richard et al. 2019) will allow us to investigate HI Ly halos for Si ii* halos detected with MUSE. BlueMUSE will also make it possible to individually detect Si ii* halos of more diverse sources and to study their statistical properties thanks to a more moderate cosmic dimming effect at (by a factor of three, compared to ). Such larger samples would be useful to improve the stacking experiments (see also Section 4.3) and to study spectral variations at different radii from the centers of galaxies as done for Ly (Wisotzki et al. 2018; Guo et al. 2024, 2023b). Such studies will be advanced more with better sensitivities and higher spatial resolutios achieved by extremely large telescopes such as Thirty-meter telescope (TMT), one of whose key science cases is CGM mapping with WFOS (Skidmore et al. 2015), as well as European-Extremely large telescope (E-ELT).